If F(x)=x-5and G(x)=x2-2x then g(f(-1)=: If F(x)=x-5and G(x)=x2-2x then g(f(-1)=
Answer by Fombitz(1755)

f(x)=x-5
g(x)=x^2-2x
First find f(-1).
f(x)=x-5
f(-1)=-1-5
f(-1)=-6
Then
g(x)=x^2-2x
g(f(-1))=g(-6)=(-6)^2-2(-6)
g(f(-1))=36+12
g(f(-1))=48